OpenStudy (raheelafzaal):

What come next in the following series 5 10 16 23 31 40 ______ a) 50 b) 49 c) 51 d) 48

OpenStudy (anonymous):

10-5= ? 16-10= ? 23-16= ? 31-23= ? 40-31= ? Fill in the blanks

OpenStudy (anonymous):

Hmm,what do u think? @raheelafzaal

OpenStudy (anonymous):

 what is the next number? @raheelafzaal